# Spectrum from Defocus: Fast Spectral Imaging with Chromatic Focal Stack

### v1

@@ 3.2 Inverse Algorithm
We solve the spectral image recovery problem with an iterative optimization that alternates between inverting the physical forward model in Eq. 1 and applying a data-driven regularizer. Specifically, we modify plug-and-play ADMM [8] with two task-specific augmentations to improve robustness and computational speed. First, we project spectra onto a limited number of principal components, taking advantage of the low-rank statistics of natural visible spectra. Second, we separate the spatial and frequency domain computations for fast matrix inversion.
@@
$\min_{\mathbf{z}}\frac{1}{2}||\mathbf{y}-\mathbf{C}\mathbf{H}\mathbf{P}\mathbf{z}||_{2}^{2}+\Phi_{\boldsymbol{\theta}}(\mathbf{P}\mathbf{z}),$ (2)
where $\Phi_{\boldsymbol{\theta}}(\cdot)$ is an off-the-shelf deep-learning-based regularizer with parameters $\boldsymbol{\theta}$. The vector $\mathbf{z}\in\mathbb{R}^{H\cdot W\cdot v}$, where $v$ represents the dimensionality of the selected eigenspace, is utilized to exploit the property that natural visible-range spectra primarily reside in a lower-dimensional space (i.e., $\mathbf{x}\approx\mathbf{P}\mathbf{z}$). We define $\mathbf{P}=\mathbf{B}^{T}\otimes\mathbf{I}_{HW}$, where the rows of $\mathbf{B}$, shaped as $\mathbb{R}^{v\times C}$, are eigenvectors obtained from the Harvard dataset [6], as reported in [2]. The matrix $\mathbf{I}_{HW}$ and the operator $\otimes$ denote an identity matrix shaped as $HW\times HW$ and the Kronecker product, respectively.
To adopt a plug-and-play ADMM [8], we transformed Eq. 2 by introducing slack variables $\mathbf{v}$ and $\mathbf{u}$, and by using the change of variables $\hat{\mathbf{H}}:=\mathbf{H}\mathbf{P}$:
@@
[…]
@@
4:     ${\color[rgb]{0,1,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,1}\pgfsys@color@cmyk@stroke{1}{0}{0}{0}\pgfsys@color@cmyk@fill{1}{0}{0}{0}\mathbf{v}_{i+1}\leftarrow(\mathbf{C}^{T}\mathbf{C}+\mu_{1}\mathbf{I})^{-1}(\mathbf{C}^{T}\mathbf{y}+\mu_{1}\hat{\mathbf{H}}\mathbf{z}_{i}-\boldsymbol{\xi}_{i})}$
@@
6:     ${\color[rgb]{0,1,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,1}\pgfsys@color@cmyk@stroke{1}{0}{0}{0}\pgfsys@color@cmyk@fill{1}{0}{0}{0}\mathbf{u}_{i+1}\leftarrow\mathbf{P}^{T}\phi_{\boldsymbol{\theta}}\left(\mathbf{P}(\mathbf{z}_{i+1}+\boldsymbol{\eta}_{i})\right)}$
7:     ${\color[rgb]{1,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,1}\pgfsys@color@cmyk@stroke{0}{1}{0}{0}\pgfsys@color@cmyk@fill{0}{1}{0}{0}\boldsymbol{\xi}_{i+1}\leftarrow\boldsymbol{\xi}_{i}+\mu_{1}(\mathbf{v}_{i+1}-\hat{\mathbf{H}}\mathbf{z}_{i+1})}$
8:     ${\color[rgb]{1,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,1}\pgfsys@color@cmyk@stroke{0}{1}{0}{0}\pgfsys@color@cmyk@fill{0}{1}{0}{0}\boldsymbol{\eta}_{i+1}\leftarrow\boldsymbol{\eta}_{i}+\mu_{2}(\mathbf{u}_{i+1}-\mathbf{z}_{i+1})}$
9:end while
Figure 3: Comparison of simulated SOTA methods. We illustrate the performance of several SOTA methods simulated on five hyperspectral images from the KAIST [9] and CAVE [30] dataset. The PSNR of the reconstructed hyperspectral image is listed for each, below an RGB projection of the data. Our method, using a simulated focal stack of 5 measurements, performs well, with high spatial resolution and no color tinting (compare to MST). Spectra are shown for the numbered points, indicated with white arrows on each image. We reconstruct spectra smoothly and accurately. Note that MST can only reconstruct a limited wavelength range, as its training data spans only $453-648$ nm.
Plug-and-play ADMM primarily consists of two main update steps: a primal variable update ($\mathbf{v}$, $\mathbf{z}$, $\mathbf{u}$) and a dual variable update ($\boldsymbol{\xi}$ and $\boldsymbol{\eta}$). In the primal update, $\mathbf{v}$ can be interpreted as a pseudo-measurement estimated from the true measurement $\mathbf{y}$ and the current update $\mathbf{z}_{i}$. An estimate of a projected image $z_{i+1}$ is then generated using a Wiener-like filter, $(\mu_{1}\hat{\mathbf{H}}^{T}\hat{\mathbf{H}}+\mu_{2}\mathbf{I})^{-1}$. Since this estimate may contain noise, the proposed algorithm further applies an off-the-shelf deep learning denoiser $\phi_{\boldsymbol{\theta}}(\cdot)$ to suppress noise and enhance the result. We applied the denoiser after converting $\mathbf{z}_{i+1}$ back to the image domain, as the off-the-shelf denoiser was trained on data in the hyperspectral image domain [18]. In the dual update, $\boldsymbol{\xi}$ and $\boldsymbol{\eta}$ are updated by evaluating the residual of the constraints in Eq. 3. By iteratively alternating between primal and dual updates, the algorithm can heuristically converge to produce a projected image, which can then be used to recover a hyperspectral image (i.e., $\mathbf{x}\approx\mathbf{P}\mathbf{z}$). Specifically, we declare convergence when the step size falls below a fixed threshold or exceeds the previous step size scaled by a convergence check parameter. It should be noted that updating $\mathbf{z}$ requires inverting a large, non-diagonal matrix. Given the matrix’s large size, computing its inverse directly is impractical. To overcome this challenge, we derived a fast matrix inversion formula by leveraging the Block Circulant with Circulant Blocks (BCCB) structure of submatrices in $\mathbf{H}$ (see Supplement). This fast inversion formula enables us to run the Plug-and-play ADMM algorithm, achieving sub-second reconstruction times. As a final step, we correct for the spectral response curve of our grayscale camera and visible light filter.
4 Experimental Results
4.1 System Specifications
We use the Thorlabs LBF254-050-A best-form lens as the objective lens and the Thorlabs AL2550M-A aspherical lens as the focusing lens, both with 50 mm focal lengths. These lenses were chosen based on their measured performance in optical simulations, with particular emphasis on their longitudinal chromatic aberration, which serves as the primary spectral encoding mechanism in our system. The system produces about 1 mm of axial chromatic focal shift across the visible spectrum, with a 34 cm depth of field ($2.64-2.98$ m). Longitudinal chromatic aberration dominates spherical by a factor of 6, and other aberrations are small. See supplement for a detailed description of the optics.
To calibrate the PSFs at different lens position and target wavelengths, we imaged a point source at $2.8$ m through a set of 12 visible bandpass filters with central wavelengths spanning $400$ to $694$ nm. Fig. 2b shows the PSF, $\mathbf{K}(z_{i},\lambda_{j})$, at several lens positions $z_{i}$ for each wavelength $\lambda_{j}$. We collect PSFs at 12 lens positions, corresponding to an in-focus image for each waveband. For each reconstruction, we use the PSFs for all 12 wavebands and only the subset of lens positions used in the measurements. We assume the target to be within the depth of field and ignore defocus in the image formation, see Fig. 5a and supplement for defocus effects.
4.2 Synthetic Experiments
We evaluated the performance of our SfD camera on there publicly available hyperspectral image datasets [9, 30, 4] and compare to several state-of-the-art methods. Photon noise was simulated for all methods using a Poisson process. Specifically, we assume a brightly lit scene with a photon flux of $4.6\cdot 10^{17}\text{m}^{-2}\text{s}^{-1}$ in the visible range.
To ensure a fair comparison across different modalities, we configured a total exposure time of 5 seconds (i.e., number of exposures $\times$ duration per exposure) for all modalities and maintained the same brightness conditions as described above. Notably, under this configuration, increasing the number of exposures does not result in a longer total exposure time. We also accounted for the blocking and absorption effects of optical components in each modality (see supplement). Additionally, we assumed an ideal camera with negligible transition time between exposures, infinite dynamic range, no quantization error, and zero readout noise.
As shown in Tab. 1, our method outperforms others in PSNR and SSIM while achieving second place in SAM. Compared to our SfD camera, tunable-filter (TF) methods and SDC exhibit lower light efficiency due to spectral filters blocking a significant portion of incoming photons. The performance of data-driven methods, such as MST and Choi et al., degrades when the noise in the training data does not align with the noise encountered during evaluation. Conversely, KRISM and Zhan et al.’s work lack spatial denoising modules, leading to inferior performance under photon-limited conditions. Noted that KRISM achieves the highest SAM performance, as it is capable to conduct singular value decomposition in the individual scene. Additionally, the algorithm used in [31] is an naive inverse filtering, which can be extremely unstable in the noisy condition.
In terms of computation and optics, our SfD camera achieves a well-balanced trade-off while delivering SOTA reconstruction performance. Although KRISM offloads computational complexity to optics, its bulky system limits applicability in portable devices. Similarly, while MST ranks second in PSNR performance, it requires approximately twice the computational time and a greater number of optical components.
As shown in Fig. 3, our SfD camera demonstrates superior reconstruction quality in low spatial-frequency regions, such as the white background of the flower and the gray background of the mug. Due to our physics-based iterative algorithm, our SfD camera exhibits minimal hallucination artifacts, as evidenced by a comparison of the yellow pepper patches in MST and our method. Notably, although Zhan et al. employ chromatic aberration in a manner similar to our approach, their inverse-filtering algorithm is highly sensitive to noise, resulting in inferior color reconstruction.
Fig. 3 also shows spectra at 6 indicated pixels, illustrating our successful high dimensional reconstructions across multiple datasets. Notably, points 5 and 6 demonstrate a metameric case between real and fake food items. Unlike the other methods, SfD is able to match both of these spectra well, providing the key material-differentiation advantage of hyperspectral imaging over conventional RGB imaging [11]. Note that the data-driven technique MST fails the most seriously on distinguishing between these spectra, as seen previously in [12] .
Figure 4: Quantitative evaluation on real data. We reconstruct the Macbeth color chart from a 5-measurement focal stack to evaluate hyperspectral and trichromatic reconstruction quality. The spectral reconstruction of the full image with three spectra from RGB patches and the $\Delta E_{00}$ perceptual error of each square are presented. The SAM in degree is also reported for the RGB color patches.
4.3 Experiments on Real Data
We built a prototype camera according to the description in Sect. 3.1, shown in Fig. 1b. We imaged and reconstructed several real scenes, with data and code publicly available on our GitHub page.
Objects were placed at a distance of $2.8$ meters from the camera, at which depth our camera has a theoretical depth of field of $34$ cm. Additionally, to evaluate light efficiency of the SfD camera, our illuminating light source delivers approximately 25 lux, which is about 15 times smaller than common indoor brightness condition [1]. We ensured that all test targets remained within the abovementioned depth of field to minimize the optical defocus caused by the depth change.
In addition to keeping the targets within the depth of field, positioning them close to the optical axis helps mitigate focal length variations caused by spherical aberration and minimizes lateral chromatic aberration. These optical effects, which are not modeled in our SfD algorithm, remain a source of model mismatch. To minimize their impact on our reconstruction algorithm, we center-crop the image around the test target, where the optical system performs closest to ideal.
Figure 5: Robustness of Color Checker Recovery. (a) We assess the reconstructed spectral accuracy at varying distances from the calibration plane. Compared to the theoretical working distance of $16$ cm from the calibration plane, mild deviations from this range do not significantly impact the results. (b) We evaluate spectral accuracy as a function of the number of measurements. Adding measurements (which in the real data setting corresponds to a longer total exposure time), improves reconstruction, with performance saturating more quickly in RGB than in spectrum. See supplement for corresponding images.
Figure 6: Qualitative results on real data. Measurements from our grayscale camera show significant blur (left). Reconstructing color from a three-measurement focal stack (middle) leads to spatial and spectral artifacts that can be resolved by including additional measurements (right).
4.3.1 Quantitative Evaluation
To assess the spectral accuracy and trichromatic color quality of our SfD algorithm, we imaged the standard Macbeth color chart, shown in Fig. 4. For spectral accuracy evaluation, we sampled a pixel from each RGB color patch in the reconstruction. We then used the Spectral Angle Mapper (SAM) to compare each sample’s spectral accuracy to the known reflectance of the patch, adjusted by the scene’s illuminance spectrum. The reconstructed image and three example spectral curves are shown in Fig. 4. The SfD camera accurately reconstructs the image and captures the overall spectral trend in the selected samples. However, its performance degrades in the longer wavelength range ($>600$ nm) due to the lower scene illuminance in this range, resulting in a reduced signal-to-noise ratio (SNR).
Additionally, we characterize the accuracy of color reproduction with respect to human vision in terms of the perceptual color difference metric, $\Delta E_{00}$[1], as follows. First, we model the known Macbeth chart reflectances $R(\lambda)$ as our ground truth reference spectra $S_{GT}(\lambda)=R(\lambda)$. We then project this ground truth spectra $S_{GT}(\lambda)$ into the CIE XYZ color space by multiplying and integrating over the human color matching functions $\bar{x},\bar{y},\bar{z}$:
$\begin{bmatrix}X\\ Y\\ Z\end{bmatrix}$ $=\begin{bmatrix}\alpha_{x}\int\bar{x}(\lambda)S(\lambda)\mathrm{d}\lambda\\ \alpha_{y}\int\bar{y}(\lambda)S(\lambda)\mathrm{d}\lambda\\ \alpha_{z}\int\bar{z}(\lambda)S(\lambda)\mathrm{d}\lambda\end{bmatrix}$ (4)
where $\alpha_{i}$ is a normalization factor. We repeat this spectra-to-XYZ process with our reconstructed spectra $S_{CfD}(\lambda)$. For the purposes of this application, where low luminance reflective surfaces are being evaluated, we use the perceptual difference metric $\Delta E_{00}$, which measures differences in the $L^{*}a^{*}b^{*}$ color space. We use the XYZ of the white patch as our adapting white point during conversion to $L^{*}a^{*}b^{*}$. To reduce the effects of non-uniform illumination in our scene and to especially highlight spectral color quality difference, the luminance between the two colors was first normalized before evaluating the color difference:
$XYZ_{CfD}=XYZ_{CfD}\cdot Y_{GT}/Y_{CfD},$ (5)
$\Delta E_{00}=(XYZ_{GT},XYZ_{CfD};XYZ_{P}).$ (6)
The $\Delta E_{00}$ values for each of the 24 Macbeth color checker squares are shown in Fig. 4. A $\Delta E_{00}$ value at or below 2 is considered perceptually equivalent. Our average $\Delta E_{00}$ color difference across the entire Macbeth chart is 3.77. Although this is above visual threshold, these differences are likely to go unnoticed in practice.
4.3.2 Qualitative Evaluation
We evaluated the robustness of the SfD camera by positioning the standard Macbeth color chart at varying distances from the calibration plane. As shown in Fig. 5a, the results indicate that the SfD camera maintains a consistent SAM level within a $30$ cm range. Compared to the theoretical working distance from the calibration plane, the SfD camera demonstrates robustness even when the theoretical assumptions are mildly violated.
Fig. 5b shows that increasing the number of measurements up to 10 improves spectral reconstruction accuracy, while RGB color improvements saturate after 5 measurements. Corresponding RGB images are shown in supplement.
Fig. 6 shows several scenes assembled in the lab, illustrating spatial and spectral effects in our reconstructions. On the left we show two example measurements. From three measurements (lens positions $z_{2}$, $z_{3}$, $z_{5}$), our reconstructed RGB images show some spatial and spectral artifacts, such as the false color patterns within the black and white pinwheel target. Including additional measurements ($z_{7}$, $z_{8}$, $z_{9}$) stabilizes our inversion and removes these false color patterns, while sharpening edges and boosting saturation.

### v2

@@ 3.2 Inverse Algorithm
We solve the spectral image recovery problem with an iterative optimization that alternates between inverting the physical forward model in Eq. 1 and applying a data-driven regularizer. Specifically, we modify plug-and-play ADMM [10] with two task-specific augmentations to improve robustness and computational speed. First, we project spectra onto a limited number of principal components, taking advantage of the low-rank statistics of natural visible spectra. Second, we separate the spatial and frequency domain computations for fast matrix inversions.
@@
[…]
where $\Phi_{\boldsymbol{\theta}}(\cdot)$ is an off-the-shelf deep-learning-based regularizer with parameters $\boldsymbol{\theta}$ [29]. The vector $\mathbf{z}\in\mathbb{R}^{H\cdot W\cdot v}$, where $v$ represents the dimensionality of the selected eigenspace, is utilized to exploit the property that natural visible-range spectra primarily reside in a lower-dimensional space (i.e., $\mathbf{x}\approx\mathbf{P}\mathbf{z}$). We define $\mathbf{P}=\mathbf{B}^{T}\otimes\mathbf{I}_{HW}$, where the rows of $\mathbf{B}$, shaped as $\mathbb{R}^{v\times C}$, are eigenvectors obtained from the Harvard dataset [8], as reported in [1]. The matrix $\mathbf{I}_{HW}$ and the operator $\otimes$ denote an identity matrix shaped as $HW\times HW$ and the Kronecker product, respectively.
To adopt a plug-and-play ADMM, we transformed Eq. 2 by introducing slack variables $\mathbf{v}$ and $\mathbf{u}$, and by using the change of variables $\hat{\mathbf{H}}:=\mathbf{H}\mathbf{P}$:
@@
$\textrm{s.t.}\quad\mathbf{v}=\hat{\mathbf{H}}\mathbf{z},~\mathbf{u}=\mathbf{z}.$
@@
4:  $\mathbf{v}_{i+1}\leftarrow(\mathbf{C}^{T}\mathbf{C}+\mu_{1}\mathbf{I})^{-1}(\mathbf{C}^{T}\mathbf{y}+\mu_{1}\hat{\mathbf{H}}\mathbf{z}_{i}-\boldsymbol{\xi}_{i})$}
5:  $\mathbf{z}_{i+1}\leftarrow(\mu_{1}\hat{\mathbf{H}}^{T}\hat{\mathbf{H}}+\mu_{2}\mathbf{I})^{-1}\Big(\hat{\mathbf{H}}^{T}(\mu_{1}\mathbf{v}_{i}+\boldsymbol{\xi}_{i})$} $\hskip 50.00008pt+(\boldsymbol{\eta}_{i}+\mu_{2}\mathbf{u}_{i})\Big)$}
6:  $\mathbf{u}_{i+1}\leftarrow\mathbf{P}^{T}\phi_{\boldsymbol{\theta}}\left(\mathbf{P}(\mathbf{z}_{i+1}+\boldsymbol{\eta}_{i})\right)$}
7:  $\boldsymbol{\xi}_{i+1}\leftarrow\boldsymbol{\xi}_{i}+\mu_{1}(\mathbf{v}_{i+1}-\hat{\mathbf{H}}\mathbf{z}_{i+1})$}
8:  $\boldsymbol{\eta}_{i+1}\leftarrow\boldsymbol{\eta}_{i}+\mu_{2}(\mathbf{u}_{i+1}-\mathbf{z}_{i+1})$}
9:end while
Figure 3: Simulated reconstruction examples. (a) Hyperspectral reconstructions from the KAIST [12] and CAVE [51] datasets, illustrated with RGB projections, annotated with hyperspectral PSNR. Insets highlight characteristic failures of each method, with comparison to ground truth. (b) RGB projections and PSNR values under a low-exposure setting (2.9 s total exposure), where SfD has an even more significant advantage (see supplement). (c) Reconstructed spectra from the white arrow locations in (a). The selected points cover a range of colors, with points 5 and 6 illustrating the importance of spectral imaging with an RGB metamer. Note that MST and Spectral DefocusCam exhibit narrower spectral coverage, due to their training/calibration data and forward models, and that tunable filters and Zhan et al. are illustrated instead in supplement due to their high spectral noise.
Plug-and-play ADMM primarily consists of two main update steps: a primal variable update ($\mathbf{v}$, $\mathbf{z}$, $\mathbf{u}$) and a dual variable update ($\boldsymbol{\xi}$ and $\boldsymbol{\eta}$). In the primal update, $\mathbf{v}$ can be interpreted as a pseudo-measurement estimated from the true measurement $\mathbf{y}$ and the current update $\mathbf{z}_{i}$. An estimate of a projected image $z_{i+1}$ is then generated using a Wiener-like filter, $(\mu_{1}\hat{\mathbf{H}}^{T}\hat{\mathbf{H}}+\mu_{2}\mathbf{I})^{-1}$. Since this estimate may contain noise, the proposed algorithm further applies an off-the-shelf deep learning denoiser [29] $\phi_{\boldsymbol{\theta}}(\cdot)$ to suppress noise and enhance the result. See supplement for a comparison demonstrating the advantage of this deep denoiser over $\ell_{1}$ regularization. We applied the denoiser after converting $\mathbf{z}_{i+1}$ back to the image domain, as the off-the-shelf denoiser was trained on data in the hyperspectral image domain. In the dual update, $\boldsymbol{\xi}$ and $\boldsymbol{\eta}$ are updated by evaluating the residual of the constraints in Eq. 3. By iteratively alternating between primal and dual updates, the algorithm can heuristically converge to produce a projected image, which can then be used to recover a hyperspectral image (i.e., $\mathbf{x}\approx\mathbf{P}\mathbf{z}$). Specifically, we declare convergence when the step size falls below a fixed threshold or exceeds the previous step size scaled by a convergence check parameter. It should be noted that updating $\mathbf{z}$ requires inverting a large, non-diagonal matrix. Given the matrix’s large size, computing its inverse directly is impractical. To overcome this challenge, we derived a fast matrix inversion formula by leveraging the Block Circulant with Circulant Blocks (BCCB) structure of submatrices in $\mathbf{H}$ (see supplement). This fast inversion formula enables us to run the full algorithm with sub-second reconstruction times. As a final step, we correct for the spectral response curve of our grayscale camera and visible light filter. Additionally, see supplement for the settings of $\mu_{1}$ and $\mu_{2}$ in simulated and real data.
4 Experimental Results
We evaluate the performance of our method against several state-of-the-art methods in simulation on three publicly available hyperspectral datasets [12, 51, 8], then demonstrate its real-world performance with a hardware prototype.
4.1 Simulation Settings
All modalities are simulated for the same total exposure time (i.e., number of exposures $\times$ duration per exposure) so that snapshot and multi-exposure methods have equivalent acquisition time and total photon budget. We assume a brightly lit condition ($7.5\cdot 10^{17}~\text{m}^{-2}\text{s}^{-1}$ in the visible range) and ideal cameras (negligible transition time between exposures and no quantization or readout noise). The simulations include Poisson noise and account for the light efficiency of optical components (see supplement). Unless otherwise stated, the total exposure time is 5 seconds. We show in Fig. 3b and supplement that the advantages of our method are even stronger for lower total exposure times.
4.2 Simulation Results
As shown in Tab. 1 and illustrated in Fig. 3, SfD achieves the best PNSR and SSIM, and second-best SAM across 9 methods. The performance of our method arises from its combination of photon-efficient optics and a physics-based reconstruction that remains stable under Poisson noise.
The spectral-filter-based modalities exhibit substantially lower spectral accuracy (see SAM in Tab. 1 and Fig. 3c1–4), primarily due to their lower photon efficiency (see supplement). The reduced photon efficiency leads to increased image noise or loss of high spatial frequency from denoising (see insets in Fig. 3), and the degradation becomes more pronounced under limited exposure conditions (see Fig. 3b).
In contrast, dispersive-based modalities somewhat mitigate photon loss at the cost of a more complex optical setup (see opt. components in Tab. 1). This optical complexity leads to relatively lower photon efficiency than SfD, causing haziness and other color errors, particularly when the photon budget is limited. Note that the purely data-driven MST method, despite its high performance in Tab. 1, exhibits severe hallucinations in the low exposure time case (see yellow/red pepper in Fig. 3b). KRISM achieves the highest SAM performance by applying per-scene singular value decomposition (SVD). However, its per-scene basis can introduce spectral bias, as the SVD optimization may favor colors specific to each scene (see oversaturation in the inset and undersaturation in the low exposure image). The DOE-based 2in1 Camera loses significant photons to its RGB filters, and its data-driven model introduces haze and undersaturation for many scenes.
Chromatic-aberration-based modalities, including Zhan et al. and ours, offer greatly improved photon efficiency (see supplement). However, high photon efficiency does not guarantee robust reconstruction. In particular, Zhan et al.’s method relies on naive inverse filtering, which becomes highly unstable even under mild noise.
Our method produces SOTA hyperspectral results, with robust performance across scenes from all datasets. Errors visible in RGB projections are rare, but see inset for color bleeding into dark regions. Small spectral errors can be seen in Fig. 3c, which shows that reconstructed spectra can be slightly oversmoothed by our use of a low-rank hyperspectral representation. Importantly, our SfD hardware preserves nearly all incident light and addresses shot noise witha robust reconstruction algorithm enables, for stable performance even under low exposure conditions. Fig. 3b shows a dramatic advantage to SfD in low light levels, demonstrating that our physics-based iterative reconstruction produces sharp, accurate results with minimal hallucination while still maintaining sub-second runtime.
Figure 4: Results from SfD prototype. (a) We show raw grayscale measurements from the camera, samples from corresponding per-channel reconstruction results, and the aligned ground truth for four real scenes. $\Delta z$ indicates the lens displacement relative to the initial position, and sampled channels have an approximate bandwidth of 1010nm centered at the displayed wavelength. The reconstructions demonstrate faithful recovery of both spatial details and spectral content. (b) Reconstructed and ground truth spectral curves for each Macbeth color patch.
4.3 Data Collection
We built a prototype according to Sec. 3. Our prototype consists of a grayscale sensor and two 50 mm off-the-shelf lenses, where the focusing lens is mounted on a linear stage (see supplement for more details including a full parts list). The two selected lenses make longitudinal chromatic aberration the dominant aberration in our system (see supplement). Our optical design has a depth of field of 34 cm (2.64–2.98 m) and an axial chromatic focal shift of approximately 0.7 mm across the visible spectrum (see supplement), which we use as the primary source of spectral encoding.
@@ 4.3 Data Collection
Ground-truth spectral images were captured using Liquid Crystal Tunable Filters (LCTFs), with the total exposure time slightly adjusted to accommodate the scene’s dynamic range. A flat spectral target was used to correct per-channel filter absorption. For all data, we consider the central $800$ by $800$ pixels (9.7 deg FOV) cropped from the $1920$ by $1200$ image. All targets were placed 2.8 m from the camera unless otherwise noted, for minimal depth defocus. See supplement for details, all code and data on our Github page.
4.4 Results on Real Data
Our SfD camera produces sharp and accurate spectral reconstructions from blurry focal stacks and exhibits high performance consistent with simulations. Fig. 4a shows a subset of the collected data and per-channel ground truth, reconstruction, and PSNR for four real scenes. Observe that a variety of colorful objects are reproduced faithfully, both within each channel and in their collective color recovery. Results on all channels for these scenes and the book in Fig. 1 are shown in supplement. For comparison to a direct measurement baseline, where chromatically-aberrated images focused at 671, 546, and 442 nm are used as RGB channels, see supplement.
Note that, while nearly all channels on nearly all scenes produce high-quality reconstructions, some channels of the pinwheel target (see 700 nm in Fig. 4a) show a failure to reconstruct the full resolution and contrast of the pattern. This is likely due to the difficulty in distinguishing between true black regions that have been brightened by defocused light from neighboring white pixels, versus gray regions. This failure case could be mitigated with a sensor of higher dynamic range, using more than five measurements, or by incorporating priors that favor target-like objects over natural scene statistics (e.g. total variation in place of or in addition to the deep denoiser).
Fig. 4b shows spectral reconstruction accuracy for all Macbeth ColorChecker patches. The results show an average PSNR of $29.54$ dB and a SAM of $7.42^{\circ}$ (compare to 30.81 dB and 7.35 in simulation from Tab. 1). Note that color checkers provide challenging input to our method due to their lack of natural texture, and that spectral errors are higher at longer wavelengths, where our camera’s focal shift is less informative (see supplement). This performance trend highlights the underlying working principle of SfD: stronger chromatic focal shifts yield better spectral separation, which in turn improves reconstruction accuracy.
We show in Fig. 5 the robustness of SfD on RGB and spectral color recovery using the color checker. Fig. 5a demonstrates that additional measurements stabilize reconstruction, with an effect that naturally saturates earlier on the lower-dimensional RGB reconstruction (see supplement for corresponding images and spectra). The RGB and hyperspectral reconstruction quality remains high within the expected depth range (white region in Fig. 5b). Outside this range, the reconstruction quality decreases steadily. The supplement shows that chromatic low-frequency artifacts only begin to appear when the object depth approaches 2.35 m—approximately 29 cm beyond the theoretical range. This demonstrates the robustness of SfD in practice, even when the working range assumption is mildly violated.
Figure 5: Robust Recovery. (a) Adding measurements improves reconstruction, with performance saturating more quickly in RGB than in spectrum. $\Delta E_{00}$ is a standard color perception metric [45], which we adopt here to evaluate RGB reconstruction quality. (b) SfD performs robustly within and somewhat beyond the predicted working range of 264-298 cm (dashed vertical line).